OpenStudy (anonymous):

After receiving a discount of 14.5% on its bulk order of toner cartridges, Johns’ Office Supply pays $4104. What was the price of the order before the discount? Round to the nearest dollar.Solve the problem by stating the equation you would use to solve and then solving it

OpenStudy (anonymous):

The discount price is equal to the original price minus the percentage discount taken off. Let OP = original price 4104 = OP - (.145)OP

OpenStudy (anonymous):

Solve the equation for OP and that is your answer.

OpenStudy (anonymous):

So I need to subtract 4104 - .145 = 4100. Is that correct?

OpenStudy (anonymous):

4104 = OP(1 - .145)

OpenStudy (anonymous):

OP = 4104/(1-.145)

OpenStudy (anonymous):

thats the final answer?

OpenStudy (anonymous):

Well, you have to compute the final answer, but yes.

OpenStudy (anonymous):

If so, I really

OpenStudy (anonymous):

If so, I am really confused

OpenStudy (anonymous):

where did you get 1 from

OpenStudy (anonymous):

Okay, let's say you want to buy something for that is $100 but the store is having a 25% off sale. You can figure out how much the item will be if you take 100 - (.25)(100) That's the original price minus the percentage off of the original price. That equals the discount price. The problem you have has you do it in reverse.

OpenStudy (anonymous):

Because the 1 is the original price.. 100% of the price.

OpenStudy (anonymous):

If you have something that is 25% off, then you are really paying 100% - 25% = 75% of the original price or 1 - .25

OpenStudy (anonymous):

So you can think of these types of problems like this. Discount Price = Original Price x (1 - Percent Off)

OpenStudy (anonymous):

do I actually subtract 1 - percent off

OpenStudy (anonymous):

Yes, you need to convert the percent to a decimal and subtract it.

OpenStudy (anonymous):

ok I get 0.855

OpenStudy (anonymous):

OP = 4104/(1-.145)

OpenStudy (anonymous):

4104 divided by .855 will be the original price.

OpenStudy (anonymous):

4800

OpenStudy (anonymous):

Okay, so now if you check your answer. 14.5% off of 4800 is 4800 - (.145)(4800) That should equal 4104 if that is correct.

OpenStudy (anonymous):

okay, thanks!